OpenStudy (anonymous):

You inherit 70,000 from a very wealthy grandparent, with the stipulation that for the 1st yr, the money must be invested in two stocks paying 4% and 10% annual interest, respectively. How much should be invested at each rate if the total interest earned for the yr is 4,000? 40,000 invested at 4% 30,000 30,000 invested at 4% 40,000 50,000 invested at 4% 20,000 20,000 invested at 4% 50,000

OpenStudy (anonymous):

how is this setup?

OpenStudy (anonymous):

this is pretty simple. lets go through all the options. what is the first option? 40000 invested at 4% and 30000 invested at 10% right? so add up the returns. what is 4% of 40000 added to 10% of 30000?

OpenStudy (anonymous):

3000+1600

OpenStudy (anonymous):

so that is 4600 total return. but the problem says that the return was only 4000. so option A is wrong. similarly work it out for the other options

OpenStudy (anonymous):

awesome, but i was looking for an algebraic solution, not a brute force. :)

OpenStudy (anonymous):

okay. lets go for the algebraic version. lets say he invests x amount of money at 4%. that means that he invests 70000-x amount of money at 10%. the total return is 4000. so can you write an expression for this?

OpenStudy (anonymous):

4000=70000-x*.1*1

OpenStudy (anonymous):

why are you multiplying 0.1 and 1? what does that accomplish?

OpenStudy (anonymous):

I=Prt

OpenStudy (anonymous):

okay, but that still accounts for only the investment of 70000-x invested at 10 percent. what about the other part?

OpenStudy (anonymous):

that is the heart of my question. how do i relate the two together?

OpenStudy (anonymous):

i = prt right? so i = (70000-x)*0.1*1 is the interest on the 70000-x. so lets call that interest i1. what is the interest on the 4% investment? call it i2.

OpenStudy (anonymous):

i2=(70000-x)*0.04*1

OpenStudy (anonymous):

i1+i2 = 4000

OpenStudy (anonymous):

no. if you invest 70000 - x at 10%, that means you invest x at 4% since x + 70000-x = 70000, right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i2=x*0.04*1

OpenStudy (anonymous):

okay that's good. excellent job. so i1 = (70000-x) * 0,1 * 1 and i2 = x * 0.04 * 1. but as you pointed out earlier, i1 + i2 = 4000. so what does that mean? can you combine everything and write down the equation?

OpenStudy (anonymous):

(70000-x) * 0.1 + 0.04x = 4000

OpenStudy (anonymous):

very good! now solve for x. keep in mind that x is the amount of money ivested at 4%

OpenStudy (anonymous):

50000 at .04 and 20000 at .1 fantastic thanks!!!

OpenStudy (anonymous):

great! good job!

OpenStudy (anonymous):

high five!